# Driving Animatronic Robot Facial Expression From Speech

### 19 Mar 2024 v1

@@ Abstract
Accurately replicating human facial expressions is crucial for natural human-computer interaction in robots [1, 2, 3]. Speech-synchronized, lifelike expressions enable social robots to achieve genuine emotional resonance with users [4, 5], underscoring the need to advance speech-driven expression generation for animatronic faces [6], heralding a paradigm shift in human-robot interactions.
@@ V-A Model Architecture
Specifically, the phoneme extractor follows the state-of-the-art transformer-based self-supervised pre-trained speech model, Wav2vec2 [31], finetuned on the CommonVoice dataset with $53$ languages and $392$ phoneme classes to enhance cross-language generalization. It outputs a phoneme logits sequence at $49\mathrm{Hz}$ followed by resampling to $25\mathrm{Hz}$ to match the robot orchestration rate. The temporal fusion module progressively fuses $K$-step neighborhood logits over a sliding window into frame-level speech embeddings, comprising $M=\log_{2}K$ stacked Resnet blocks with a kernel size of $3$, stride of $2$, and $H$ filters. The $N$-way speaking style vector accounts for cross-subject variations and is embedded with a fully-connected (FC) layer having $H$ hidden units. The LBS encoder takes the late-fused conditioned speech embedding as input and outputs blendshape coefficients with two successive layers: an FC layer having $2H$ hidden units with a Tanh activation, and an FC layer with $H$ hidden units followed by a Sigmoid activation.
@@ V-D Implementation Details
The model is implemented in PyTorch. The sliding window size $K$ is $8$ ($320\mathrm{ms}$). The hidden dimension $H$ is $64$. The robot has $U=4792$ vertices and $B=51$ blendshapes. Training uses the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$, learning rate $1e^{-4}$, weight decay $1e^{-4}$, and batch size $128$. $w_{m}$ is set to $1$. A dropout layer with a dropout rate of $0.1$ is added after each FC layer except the one with Sigmoid activation. $100$ epochs are trained on an NVIDIA 4090 GPU ($\sim 15\mathrm{sec}/\mathrm{epoch}$). Wav2vec2-XLSR53 is fixed during training. For inference, an $8\mathrm{Hz}$ $5$th-order low-pass Butterworth filter is applied. The model generates blendshape coefficients from speech at $>4000\mathrm{Hz}$ on GPU.

### 21 Mar 2024 v2

@@ Abstract
Accurately replicating human facial expressions is crucial for natural human-robot interaction [1, 2, 3]. Speech-synchronized, lifelike expressions enable social robots to achieve genuine emotional resonance with users [4, 5], underscoring the need to advance speech-driven expression generation for animatronic faces [6], heralding a paradigm shift in human-robot interactions.
@@ V-A Model Architecture
Specifically, the phoneme extractor follows the state-of-the-art transformer-based self-supervised pre-trained speech model, Wav2vec2 [31], finetuned on the CommonVoice dataset with $53$ languages and $392$ phoneme classes to enhance cross-language generalization. It outputs a phoneme logits sequence at $49\mathrm{Hz}$ followed by resampling to $25\mathrm{Hz}$ to match the robot orchestration rate. The temporal fusion module progressively fuses $K$-step neighborhood logits over a sliding window into frame-level speech embeddings, comprising $M=\log_{2}K$ stacked Resnet blocks with a kernel size of $3$, stride of $2$, and $H$ filters. The $N$-way speaking style vector accounts for cross-subject variations and is embedded with a fully-connected (FC) layer having $H$ hidden units. The LBS encoder takes the late-fused conditioned speech embedding as input and outputs blendshape coefficients with two successive layers: an FC layer having $2H$ hidden units with a ReLU activation, and an FC layer with $H$ hidden units followed by a Sigmoid activation.
@@ V-D Implementation Details
The model is implemented in PyTorch. The sliding window size $K$ is $8$ ($320\mathrm{ms}$). The hidden dimension $H$ is $64$. The robot has $U=4792$ vertices and $B=51$ blendshapes. Training uses the Adam optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.99$, learning rate $1e^{-4}$, weight decay $1e^{-4}$. $w_{m}$ is set to $1$. A dropout layer with a dropout rate of $0.1$ is added after each FC layer except the one with Sigmoid activation. $200$ epochs are trained on an NVIDIA 4090 GPU ($\sim 3\mathrm{sec}/\mathrm{epoch}$). Wav2vec2-XLSR53 is fixed during training. For inference, an $7\mathrm{Hz}$ $5$th-order low-pass Butterworth filter is applied. The model generates blendshape coefficients from speech at $>4000\mathrm{Hz}$ on GPU.